Let consider a company that produces TV. The mean lifetime of a TV is 6 years and
the standard deviation of the lifetime is 2.5 years. Consider a random sample of 50
TVs. Approximate the probability that the mean lifetime of these 50 TVs will exceed 7
years.
A. 0.0023 B. 0.9977 C. 0.6554 D. 0.3446 E. none of the preceding
Answer #1

Solution :

\sigma\bar x = \sigma / \sqrt n = 2.5 / \sqrt 50 = 0.3536

P(\bar x > 7) = 1 - P(\bar x < 7)

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (7 - 6) / 0.3536]

= 1 - P(z < 2.83)

= 1 - 0.9977

= 0.0023

Probability = 0.0023

option A.